OpenStudy (anonymous):

which equation is true? A. 25-(-18)=7 B. 1over3(-9)-6=-9 C. 25(-2)+7=-39+4 D. -19+8(-2)=-7(-5)

OpenStudy (anonymous):

b

OpenStudy (anonymous):

1/3. -9 = -3-6=-9

OpenStudy (anonymous):

-9=-9